### v1

@@ 2. Notations
We let $d\geq 1$ be an integer. We write $G=\operatorname{SL}_{d+1}(\mathbb{R})$, fix $\Lambda\subseteq G$ to be a lattice, and set $X=G/\Lambda$ the quotient space. We let $m_{G}$ denote the $G$-invariant Borel measure on $G$ normalised so that the induced finite measure $m_{X}$ on $X$ has tatal mass $1$. Both $m_{G}$ and $m_{X}$ are referred to as Haar measures.
@@
In the Euclidean space $\mathbb{R}^{d}$, we set $B^{\mathbb{R}^{d}}_{\rho}$ the open ball of radius $\rho$ centered at the origin. On some rare occasions (Section˜3), we might use $B_{\rho}$ as a shorthand for $B^{\mathbb{R}^{d}}_{\rho}$. We will explicitely warn about this exception at the few places it occurs.
@@
$\phi(\boldsymbol{s})=r_{\phi}O_{\phi}\boldsymbol{s}+\mathtt{b}_{\phi},\quad\boldsymbol{s}\in\mathbb{R}^{d},$
for some $O_{\phi}\in\operatorname{SO}_{d}(\mathbb{R})$, $r_{\phi}>0$ and $\mathtt{b}_{\phi}\in\mathbb{R}^{d}$.
@@
$A^{\prime}=\{\,a(t):=\operatorname{diag}(t^{\frac{1}{d+1}},\cdots,t^{\frac{1}{d+1}},t^{-\frac{d}{d+1}}):t\in\mathbb{R}_{>0}\,\}$
@@
$\int_{P^{\prime}}|\mathtt{r}_{g}|^{\varepsilon}+|\mathtt{r}_{g}^{-1}|^{\varepsilon}+\lVert\mathtt{b}_{g}\rVert^{\varepsilon}d\mu(g)<\infty.$
We assume that $\lambda$ is strongly irreducible. This means that for every set $E\subseteq\mathbb{R}^{d}$ which is a finite union of affine subspaces and satisfies $\phi E=E$ for all $\phi\in\operatorname{supp}\lambda$, we have $E=\mathbb{R}^{d}$.
@@
(2.1) $\ell=-\int_{P^{\prime}}\log\mathtt{r}_{g}\,\mathrm{d}\mu(g)$
@@
By a theorem of Bougerol-Picard [12, Theorem 2.5], the existence of a $\lambda$-stationary probability measure is equivalent to the condition $\ell>0$, i.e. the random walk on $\mathbb{R}^{d}$ driven by $\lambda$ is contractive in average. Moreover, in this case, the $\lambda$-stationary probability measure is unique, see [12, Corollary 2.7].
Finite time approximation. For any $n\in\mathbb{N}$, let $\sigma^{(n)}=\lambda^{*n}*\delta_{\boldsymbol{0}}$ be a probability measure on $\mathbb{R}^{d}$, where $\delta_{\boldsymbol{0}}$ is the Dirac measure at $\boldsymbol{0}\in\mathbb{R}^{d}$. Note that $\sigma^{(n)}$ is the image measure of $\mu^{*n}$ under the map $g\in P^{\prime}\mapsto\mathtt{b}_{g}\in\mathbb{R}^{d}$. It is known that $\sigma^{(n)}$ converges to $\sigma$ exponentially fast. More precisely, denote by $\operatorname{Lip}(\mathbb{R}^{d})$ the space of bounded Lipschitz functions on $\mathbb{R}^{d}$ equipped with the Lipschitz norm:
$\lVert f\rVert_{\operatorname{Lip}}=\lVert f\rVert_{\infty}+\sup_{\boldsymbol{s_{1}}\neq\boldsymbol{s}_{2}}\frac{|f(\boldsymbol{s}_{1})-f(\boldsymbol{s}_{2})|}{\lVert\boldsymbol{s}_{1}-\boldsymbol{s}_{2}\rVert}.$
We then have by [7, Lemma 2.2] [^3] The proof of [7] is formulated for $d=1$, but it easily carries to arbitrary $d$ by using Lemma 3.1 below in the place of [7, Lemma 2.1 (i)].:
Lemma 2.1 ([7]).
There exist constants $C,\varepsilon>0$ such that for all $n\geq 0$, $f\in\operatorname{Lip}(\mathbb{R}^{d})$, we have
$|\sigma^{(n)}(f)-\sigma(f)|\leq Ce^{-\varepsilon n}\|f\|_{\operatorname{Lip}}.$
Intervals. For real numbers $a<b$, we write $\llbracket a,b\rrbracket$ to denote $\mathbb{Z}\cap[a,b]$. Similarly, we set $\rrbracket a,b\rrbracket:=\mathbb{Z}\cap(a,b]$. We also write $\mathbb{N}_{\geq a}:=\mathbb{N}\cap[a,+\infty)$.
@@
3. Regularity of self-similar measures
We recall that $\sigma$ has a finite moment of positive order. We then establish that $\sigma$ cannot be concentrated near proper affine subspaces, and more generally near proper algebraic subvarieties of $\mathbb{R}^{d}$.
We start with the control of the tail probabilities. It can be seen as a non-concentration property near infinity.
@@ 3. Regularity of self-similar measures
$\int_{\mathbb{R}^{d}}\lVert s\rVert^{\gamma}\,\mathrm{d}\sigma(s)<\infty.$
@@
As $\lambda$ has finite exponential moment and is contractive in average, this follows from Kloeckner [39, Theorem 3.1 & Lemma 3.9]. ∎
@@
We now provide a Hölder control on the mass granted by $\sigma$ to neighborhoods of affine subspaces.
Proposition 3.2 (Non-concentration near affine subspaces).
@@
In the case where $\sigma$ arises from a self-similar IFS which is finite, contractive, and satisfies the open set condition, the result is a consequence of [33, Lemmas 8.2, 8.3]. In [33], the argument is roughly as follows. Assume for simplicity $\lambda$ has equal contraction ratios $\varepsilon\in(0,1)$. Consider $\underline{\phi}=(\phi_{i})_{i\geq 1}\sim\lambda^{\otimes\mathbb{N}}$ and observe $x_{{\underline{\phi}}}:=\lim_{k}\phi_{1}\dots\phi_{k}(0)$ has law $\sigma$. As $\|x_{{\underline{\phi}}}-\phi_{1}\dots\phi_{n}(0)\|\ll\varepsilon^{n}$, the inclusion $x_{{\underline{\phi}}}\in{\mathcal{L}}^{(\varepsilon^{n})}$ implies for every $i\leq n$ that $\phi_{1}\dots\phi_{i}(0)\in{\mathcal{L}}^{(O(\varepsilon^{i}))}$, i.e. $\phi_{i}(0)\in(\phi_{1}\dots\phi_{i-1})^{-1}{\mathcal{L}}^{(O(\varepsilon))}$. Up to taking $\varepsilon$ small enough (by considering $\mu^{*n}$ instead of $\mu$), the probability of the latter event conditionally to previous steps is smaller than $1/2$, whence the desired decay.
In our situation, the support of $\sigma$ may be unbounded, hence knowing that $x_{{\underline{\phi}}}$ is close to $\mathcal{L}$} does not mean that all steps leading to $x_{{\underline{\phi}}}$ will be as well. To deal with this difficulty, we propose an induction scheme that both takes into the count the position of $\phi_{1}\dots\phi_{n}(0)$ with respect to $\mathcal{L}$} but also how far $\mathcal{L}$} is from the origin.
@@
Given $s\in(0,1)$ and $T\in\mathbb{R}_{\geq 0}$, we set
$I_{s,T}:=\sup\{\sigma(L^{(s)})\,:\,{\mathcal{L}}\subsetneq\mathbb{R}^{d}\text{ affine with }\operatorname{dist}(0,L)\geq T\}.$
We will show that for some $C,c>0$ depending on $\sigma$ only, for all $s\in(0,1)$, $T\in\mathbb{R}_{\geq 0}$, we have
$I_{s,T}\leq Cs^{c}(1+T)^{-c}.$
Taking $T=0$, we obtain in particular Proposition˜3.2. (In fact there is an equivalence, using that $\sigma$ has finite positive moment - Lemma˜3.1 - and interpolation).
We exhibit some relations between the terms $I_{s,T}$ that will allow to perform an inductive argument. We consider parameters $\varepsilon,c\in(0,1)$ to be chosen later depending on $\sigma$. We aim to show a bound of the form:
(3.1) $\forall s\geq\varepsilon^{k},\,T\geq 0,\,\,\,\,\,\,\,I_{s,T}\leq C_{k}s^{c}(1+T)^{-c}.$

### v2

@@ 2. Notations
We let $d\geq 1$ be an integer. We write $G=\operatorname{SL}_{d+1}(\mathbb{R})$, fix $\Lambda\subseteq G$ to be a lattice, and set $X=G/\Lambda$ the quotient space. We let $m_{G}$ denote the $G$-invariant Borel measure on $G$ normalised so that the induced finite measure $m_{X}$ on $X$ has total mass $1$. Both $m_{G}$ and $m_{X}$ are referred to as Haar measures.
@@
In the Euclidean space $\mathbb{R}^{d}$, we set $B^{\mathbb{R}^{d}}_{\rho}$ the open ball of radius $\rho$ centered at the origin. On some rare occasions (Section˜4), we might use $B_{\rho}$ as a shorthand for $B^{\mathbb{R}^{d}}_{\rho}$. We will explicitely warn about this exception at the few places it occurs.
@@
$\phi(\boldsymbol{s})=\mathtt{r}_{\phi}O_{\phi}\boldsymbol{s}+\mathtt{b}_{\phi},\qquad\boldsymbol{s}\in\mathbb{R}^{d},$
for some $O_{\phi}\in\operatorname{SO}_{d}(\mathbb{R})$, $\mathtt{r}_{\phi}>0$ and $\mathtt{b}_{\phi}\in\mathbb{R}^{d}$.
@@
$A^{\prime}=\{\,a(t):=\operatorname{diag}(t^{\frac{1}{d+1}},\dots,t^{\frac{1}{d+1}},t^{-\frac{d}{d+1}}):t\in\mathbb{R}_{>0}\,\}$
@@
$\int_{P^{\prime}}\mathtt{r}_{g}^{\varepsilon}+\mathtt{r}_{g}^{-\varepsilon}+\lVert\mathtt{b}_{g}\rVert^{\varepsilon}\,\mathrm{d}\mu(g)<\infty.$
We assume that $\lambda$ is irreducible. This means that for every affine subspace $E\subseteq\mathbb{R}^{d}$ which satisfies $\phi E=E$ for all $\phi\in\operatorname{supp}\lambda$, we have $E=\mathbb{R}^{d}$.
@@
$\ell=-\int_{P^{\prime}}\log\mathtt{r}_{g}\,\mathrm{d}\mu(g)$ (2.1)
@@
By a theorem of Bougerol-Picard [12, Theorem 2.5], the existence of a $\lambda$-stationary probability measure is equivalent to the condition $\ell>0$, i.e., the random walk on $\mathbb{R}^{d}$ driven by $\lambda$ is contractive in average. Moreover, in this case, the $\lambda$-stationary probability measure is unique, see [12, Corollary 2.7].
Intervals. We write $\mathbb{N}^{*}$ for the set of positive integers. For real numbers $a<b$, we write $\llbracket a,b\rrbracket$ to denote $\mathbb{Z}\cap[a,b]$. Similarly, we set $\rrbracket a,b\rrbracket:=\mathbb{Z}\cap(a,b]$.
@@
3. Moment and finite approximation of self-similar measures
We record standard facts about $\sigma$ and how it is approximated by the translation component $\mathtt{b}_{\phi}$ of $\phi\sim\lambda^{n}$ when $n$ is large.
@@ 3. Moment and finite approximation of self-similar measures
$\int_{\mathbb{R}^{d}}\lVert\boldsymbol{s}\rVert^{\gamma}\,\mathrm{d}\sigma(\boldsymbol{s})<+\infty.$
@@
As $\lambda$ has finite exponential moment and is contractive in average, this follows from Kloeckner [46, Theorem 3.1 & Lemma 3.9]. ∎
By Bougerol-Picard [12, Theorem 2.4], there exists a measurable map $(\operatorname{Sim}(\mathbb{R}^{d}))^{\mathbb{N}^{*}}\rightarrow\mathbb{R}^{d},{\underline{\phi}}\mapsto\mathtt{b}_{{\underline{\phi}}}$ such that for $\lambda^{\otimes\mathbb{N}^{*}}$-almost every ${\underline{\phi}}=(\phi_{i})_{i\geq 1}$, one has
$\mathtt{b}_{{\underline{\phi}}}=\lim_{n\to+\infty}\mathtt{b}_{\phi_{1}\dotsm\phi_{n}}.$
Moreover, the variable $(\mathtt{b}_{{\underline{\phi}}})_{{\underline{\phi}}\sim\lambda^{\otimes\mathbb{N}^{*}}}$ has law $\sigma$. The next lemma quantifies the rate of approximation of $\mathtt{b}_{{\underline{\phi}}}$ by $\mathtt{b}_{\phi_{1}\dotsm\phi_{n}}$.
Lemma 3.2 (Finite approximation).
There exists $\gamma>0$ such that for every $p\geq 1$,
$\lambda^{\otimes\mathbb{N}^{*}}\{\,{\underline{\phi}}\,:\,\exists n\geq p,\,\|\mathtt{b}_{\phi_{1}\dotsm\phi_{n}}-\mathtt{b}_{{\underline{\phi}}}\|\geq e^{-\gamma n}\,\}\ll e^{-\gamma p}.$
Proof.
It is enough to argue for $p=n$, in other words to show that
$\lambda^{\otimes\mathbb{N}^{*}}\{\,{\underline{\phi}}\,:\,\|\mathtt{b}_{\phi_{1}\dotsm\phi_{n}}-\mathtt{b}_{{\underline{\phi}}}\|\geq e^{-\gamma n}\,\}\ll e^{-\gamma n}.$
Note that $\mathtt{b}_{\phi_{1}\dotsm\phi_{n}}=\phi_{1}\dotsm\phi_{n}(0)$ and $\mathtt{b}_{{\underline{\phi}}}=\phi_{1}\dotsm\phi_{n}(\mathtt{b}_{{\underline{\phi}}_{>n}})$ where ${\underline{\phi}}_{>n}:=(\phi_{i})_{i\geq n+1}$. In particular, for ${\underline{\phi}}\sim\lambda^{\otimes\mathbb{N}^{*}}$, the variable $\mathtt{b}_{{\underline{\phi}}_{>n}}$ has law $\sigma$ and is independent from $\phi_{1},\dotsc,\phi_{n}$. We deduce
$\lambda^{\otimes\mathbb{N}^{*}}\{\,{\underline{\phi}}\,:\,\lVert\mathtt{b}_{\phi_{1}\dotsm\phi_{n}}-\mathtt{b}_{{\underline{\phi}}}\rVert\geq e^{-\gamma n}\,\}\leq(\lambda^{*n}\otimes\sigma)\{\,(\phi,\boldsymbol{s})\,:\,\lVert\mathtt{r}_{\phi}\boldsymbol{s}\rVert\geq e^{-\gamma n}\,\}.$
The result then follows from the large deviation principle for $(\log\mathtt{r}_{\phi})_{\phi\sim\lambda^{*n}}$ and the moment control on $\sigma$ from Lemma˜3.1. ∎
As a consequence, we obtain a joint moment estimate along a $\lambda$-trajectory.
Lemma 3.3 (Moment estimate on right trajectories).
There exists $\gamma>0$ such that for every $R>1$,
$\lambda^{\otimes\mathbb{N}^{*}}\{\,{\underline{\phi}}\,:\,\exists n\geq 1,\,\lVert\mathtt{b}_{\phi_{1}\dotsm\phi_{n}}\rVert\geq R\,\}\ll R^{-\gamma}.$
Proof.
For $\eta\in(0,1)$ and $0<\gamma\lll\eta$, the domination
$\lambda^{\otimes\mathbb{N}^{*}}\{\,{\underline{\phi}}\,:\,\exists n\geq 1+\eta\log R,\,\lVert\mathtt{b}_{\phi_{1}\dotsm\phi_{n}}\rVert\geq R\,\}\ll R^{-\gamma}$
is a direct consequence of Lemmas 3.1, 3.2. It remains to show that choosing $\eta\lll 1$, we have
$\lambda^{\otimes\mathbb{N}^{*}}\{\,{\underline{\phi}}\,:\,\exists n\leq 1+\eta\log R,\,\lVert\mathtt{b}_{\phi_{1}\dotsm\phi_{n}}\rVert\geq R\,\}\ll R^{-\gamma}.$ (3.1)
For this, observe $\mathtt{b}_{\phi_{1}\dotsm\phi_{n}}=\sum_{i=0}^{n-1}\mathtt{r}_{\phi_{1}}\dotsm\mathtt{r}_{\phi_{i}}O_{\phi_{1}}\cdots O_{\phi_{i}}\mathtt{b}_{\phi_{i+1}}$. Using the Markov inequality, we may control the norm of each summand: for every $\varepsilon>0$,
$\lambda^{\otimes\mathbb{N}^{*}}\{\,{\underline{\phi}}\,:\,\mathtt{r}_{\phi_{1}}\dotsm\mathtt{r}_{\phi_{i}}\lVert\mathtt{b}_{\phi_{i+1}}\rVert\geq R^{1/2}\,\}\leq R^{-\varepsilon/2}\mathbb{E}(\mathtt{r}_{\phi}^{\varepsilon})^{i}\mathbb{E}(\lVert\mathtt{b}_{\phi_{i+1}}\rVert^{\varepsilon}).$
Fixing $\varepsilon=\varepsilon(\lambda)>0$ so that the right-hand side is finite, then choosing $\eta\lll\varepsilon$, the right hand side is smaller than $R^{-\varepsilon/4}$, and summing over $i\in\llbracket 1,1+\eta\log R\rrbracket$, Equation (3.1) follows. This concludes the proof. ∎
4. Regularity of self-similar measures
The main goal of this section to is to establish ˜4.4 stating that the self-similar measure $\sigma$ cannot be concentrated near proper algebraic subvarieties of $\mathbb{R}^{d}$.
The property that self-similar measures or sets avoid submanifolds dates back to [32, 51], then [33, 27] in the context of self-conformal or self-affine measures. Those results however all care about the mass given by $\sigma$ to a manifold, and ignore neighborhoods. Non-concentration estimates on neighborhoods can be found in [43, Theorem 7.3] in the context of absolutely friendly measures. These comprise Hausdorff measures on self-similar sets induced by finite contractive IFS’s satisfying the open set condition [43, Theorem 2.3], but not the generality allowed in this article. In the very recent paper [30], the case of Gibbs measures on self-conformal sets is investigated for neighborhoods of spheres [30, Theorem 6.1]. The present section complements these works by dealing with arbitrary self-similar measures (coming from a randomized irreducible self-similar IFS with finite exponential moment) and arbitrary subvarieties.
4.1. Hölder regularity near affine subspaces
In the next lemma, we first observe that $\sigma$ cannot charge any proper affine subspace.
Lemma 4.1.
For every proper affine subspace ${\mathcal{L}}\subsetneq\mathbb{R}^{d}$, we have $\sigma({\mathcal{L}})=0$. Equivalently, any set $E\subseteq\mathbb{R}^{d}$ which is a finite union of affine subspaces and is invariant under $\operatorname{supp}\lambda$ must satisfy $E=\mathbb{R}^{d}$.
The second characterization is a property of strong irreducibility. It strengthens the irreducibility assumption on $\lambda$ introduced in Section˜2.
Proof.
We start by establishing the equivalence.
$(\implies)$ Assume there exists $E\subsetneq\mathbb{R}^{d}$ which is a finite union of affine subspaces and $\operatorname{supp}\lambda$-invariant. Given any $x\in E$, we have $\lambda^{*n}*\delta_{x}\rightharpoonup^{*}\sigma$. Therefore $\sigma(E)=1$ by invariance, and it follows $\sigma$ gives positive mass to one of the proper affine subspaces making $E$.
[…]
Now that we have established the equivalence, we show strong irreducibility. Let $E$ be as in the statement. It can be written in a unique way as $E=\bigcup_{i\in I}E_{i}$, where $I$ is finite and $E_{i}$ are affine subspaces satisfying $E_{i}\not\subseteq E_{j}$ for all $i\neq j\in I$. To show $E=\mathbb{R}^{d}$, we may assume that $\operatorname{supp}\lambda$ permutes transitively the $E_{i}$’s and hence they all have same dimension, say $k$. We may also assume that $k$ is minimal among all possible choices of $E$. The minimality implies the $E_{i}$’s are mutually disjoint. We show that in fact $|I|=1$, i.e., there is only one component $E_{i}$. Indeed, otherwise, the quantity $\delta:=\inf\{\,\lVert x-y\rVert\,:\,{i\neq j},\,x\in E_{i},\,y\in E_{j}\,\}$ would be strictly positive. But recall that $\lambda$ is contractive in average, whence there is some similitude $\phi\in\operatorname{supp}\lambda$ with contraction ratio $\mathtt{r}_{\phi}\in(0,1)$. Then $\delta$ gets strictly smaller when we replace each $E_{i}$ by $\phi E_{i}$, thus contradicting invariance of the union. Therefore $|I|=1$. By irreducibility of $\lambda$, we deduce $E=\mathbb{R}^{d}$. ∎
@@ 4.1. Hölder regularity near affine subspaces
We upgrade Lemma˜4.1 to a Hölder control on the mass granted by $\sigma$ to neighborhoods of affine subspaces.
Proposition 4.2 (Hölder regularity near affine subspaces).
@@
In the case where $\sigma$ arises from a self-similar IFS which is finite, contractive, and satisfies the open set condition, the result is a consequence of [43, Lemmas 8.2, 8.3]. In [43], the argument is roughly as follows. Assume for simplicity $\lambda$ has equal contraction ratios $\varepsilon\in(0,1)$. Consider $\underline{\phi}=(\phi_{i})_{i\geq 1}\sim\lambda^{\otimes\mathbb{N}}$ and observe $x_{{\underline{\phi}}}:=\lim_{k}\phi_{1}\dotsm\phi_{k}(0)$ has law $\sigma$. As $\|x_{{\underline{\phi}}}-\phi_{1}\dotsm\phi_{n}(0)\|\ll\varepsilon^{n}$, the inclusion $x_{{\underline{\phi}}}\in{\mathcal{L}}^{(\varepsilon^{n})}$ implies for every $i\leq n$ that $\phi_{1}\dotsm\phi_{i}(0)\in{\mathcal{L}}^{(O(\varepsilon^{i}))}$, i.e., $\phi_{i}(0)\in(\phi_{1}\dotsm\phi_{i-1})^{-1}{\mathcal{L}}^{(O(\varepsilon^{i}))}$. Up to taking $\varepsilon$ small enough (by considering $\lambda^{*n}$ instead of $\lambda$), the probability of the latter event conditionally to previous steps is smaller than $1/2$, whence the desired decay.
In our situation, the support of $\sigma$ may be unbounded, hence knowing that $x_{{\underline{\phi}}}$ is close to $\mathcal{L}$} does not mean that all steps leading to $x_{{\underline{\phi}}}$ will be as well. To deal with this difficulty, we will use an induction scheme that both takes into account the position of $\phi_{1}\dotsm\phi_{n}(0)$ with respect to $\mathcal{L}$} but also how far $\mathcal{L}$} is from the origin.
As a preparation for the proof, we first introduce a stopping time which will be helpful in the case where the contraction ratio $\mathtt{r}_{\phi}$ is not uniform for $\phi\in\operatorname{supp}\lambda$. Given $\varepsilon>0$, we define $\tau_{\varepsilon}:\operatorname{Sim}(\mathbb{R}^{d})^{\mathbb{N}^{*}}\rightarrow\mathbb{N}\cup\{\infty\}$ to be the stopping time
$\tau_{\varepsilon}({\underline{\phi}})=\inf\{n\geq 1\,:\,\mathtt{r}_{\phi_{1}\dotsm\phi_{n}}<\varepsilon\}.$
where $\phi_{i}$ denotes the $i$-th coordinate of $\underline{\phi}$}, i.e. ${\underline{\phi}}=(\phi_{i})_{i\geq 1}$. As $\lambda$ is contracting in average, we have that $\tau_{\varepsilon}$ is $\lambda^{\otimes\mathbb{N}^{*}}$-almost surely finite. Denote by $\lambda^{*\tau_{\varepsilon}}$ the distribution of $\phi_{1}\dotsm\phi_{\tau_{\varepsilon}({\underline{\phi}})}$ as ${\underline{\phi}}\sim\lambda^{\otimes\mathbb{N}^{*}}$. The following lemma summarises what we need about this measure.
Lemma 4.3.
For all $\delta,\gamma>0$ small enough depending on $\lambda$, for all $\varepsilon\in(0,1)$, we have the following.
  1. (1)
    The measure $\sigma$ is $\lambda^{*\tau_{\varepsilon}}$-stationary.
  2. (2)
    The variable $\varepsilon/\mathtt{r}_{\phi}$ where $\phi\sim\lambda^{*\tau_{\varepsilon}}$ has a moment of order $\gamma$ bounded by $2$:
    $\int(\varepsilon/\mathtt{r}_{\phi})^{\gamma}\,\mathrm{d}\lambda^{*\tau_{\varepsilon}}(\phi)\leq 2.$ (4.1)
  3. (3)
    Provided that $\varepsilon\in(0,\delta)$, for all proper affine subspaces ${\mathcal{L}}\subseteq\mathbb{R}^{d}$,
    $\lambda^{*\tau_{\varepsilon}}\{\,\phi\,:\,\operatorname{dist}(\mathtt{b}_{\phi},{\mathcal{L}})\leq\delta\,\}\leq 2^{-9}.$ (4.2)
  4. (4)
    For all $T\geq\delta^{-1}/2$,
    $\lambda^{*\tau_{\varepsilon}}\{\,\phi\,:\,\lVert\mathtt{b}_{\phi}\rVert\geq T\,\}\leq 2^{-9}(1+T)^{-\gamma}.$ (4.3)
@@
  1. (1)
    This is a consequence of [6, Lemma A.2].
  2. (2)
    By [7, Lemma A.19], we have for every $C>1$,
    $\lambda^{*\tau_{\varepsilon}}\{\mathtt{r}_{\phi}<\varepsilon 2^{-C}\}\leq\sum_{n=0}^{+\infty}\lambda^{*\tau_{1/2}}\{\mathtt{r}_{\phi}<2^{-n-C}\},$
    and it follows that
    $\lambda^{*\tau_{\varepsilon}}\{\mathtt{r}_{\phi}<\varepsilon 2^{-C}\}\leq\int_{|\log\mathtt{r}_{\phi}|\geq C\log 2}1+|\log\mathtt{r}_{\phi}|\,\mathrm{d}\lambda^{*\tau_{1/2}}(\phi)$
    This integral is bounded by $O(e^{-\alpha C})$ for $\alpha\lll 1$, because finite exponential moment holds for $\lambda$, whence for $\tau_{1/2}$ (by the large deviation principle), and then for $\lambda^{*\tau_{1/2}}$. This justifies (4.1).
  3. (3)
    In view of Lemma˜4.1 and using a compactness argument, we can find $\eta=\eta(\sigma)>0$ such that $\sup_{{\mathcal{L}}}\sigma({\mathcal{L}}^{(2\eta)})\leq 2^{-10}$. On the other hand, Lemma˜3.2 implies that for $\delta>0$ small enough depending on $\lambda,\eta$, for $\varepsilon\in(0,\delta)$,
    $\lambda^{\otimes\mathbb{N}^{*}}\{{\underline{\phi}}\,:\,\lVert\mathtt{b}_{\phi_{1}\dotsm\phi_{\tau_{\varepsilon}({\underline{\phi}})}}-\mathtt{b}_{\underline{\phi}}\rVert\geq\eta\}\leq 2^{-10}.$
    The claim (4.2) follows by combining the two inequalities.
  4. (4)
    It is a direct consequence of Lemma˜3.3 and the Markov inequality. ∎
We may now pass to the proof of Proposition˜4.2.
Proof of Proposition˜4.2.
Let $c,\varepsilon\in(0,1)$ be some parameters that we will specify below. Provided $c$ is sufficiently small depending on $\lambda$, and $\varepsilon$ is well chosen [^4] It will be important to $\varepsilon$ be fine tuned depending on $c$, in particular not too small and not too large. depending on $c$, we show the following statement (which clearly implies the desired upper bound): For every $k\in\mathbb{N}$, for all proper affine subspaces ${\mathcal{L}}\subsetneq\mathbb{R}^{d}$, all $s\geq\varepsilon^{k}$,
$\sigma\bigl({\mathcal{L}}^{(s)}\bigr)\leq C_{k}s^{c}\bigl(1+\operatorname{dist}(0,{\mathcal{L}})\bigr)^{-c},$ (4.4)
where $C_{k}=\varepsilon^{-kc/2}C_{0}$ for $k\geq 1$ and $C_{0}=C_{0}(\sigma)$ depends only on $\sigma$.